OpenStudy (anonymous):

n/3+n/4=1/2 Can anybody solve that and explain it to me?

OpenStudy (vijay):

[ 4n + 3n] / 12 = 1/2 7n = 6 n = 6/7

OpenStudy (anonymous):

make your life easy by multiplying both sides by 12 get \[12\times (\frac{n}{3}+\frac{n}{4})=\frac{1}{2}\times 12\] \[4n+3n=6\] \[7n=6\] \[n=\frac{6}{7}\]